### v1

@@ Abstract
Self-driving labs (SDLs), employing automation and machine learning (ML) to accelerate experimental procedures, have enormous potential in the discovery and optimisation of new materials. However, in thin film science, SDLs are mainly restricted to solution-based synthetic methods which are easier to automate but cannot access much of the broad chemical space of inorganic materials. This work focuses on a developing SDL based on magnetron co-sputtering. We focus here on a key challenge for high-throughput materials exploration using combinatorial frameworks, namely obtaining accurate composition maps on multi-element, compositionally graded thin films. This normally requires time-consuming ex-situ analysis prone to systematic errors. We present an in-situ, machine-learning driven approach to produce composition maps for arbitrary source combinations and sputtering conditions, that is both rapid and calibration-free. We develop a method to accurately and automatically learn and predict the composition distribution in a multi-element combinatorial thin film, using in-situ measurements from quartz-crystal microbalance sensors at three locations in a sputter chamber. For a given sputtering source, the sensor readings are sequentially learned as a function of the sputtering pressure and magnetron power, through active learning using Gaussian processes (GPs). The final GPs are thereafter combined with a geometric model of the deposition flux distribution in the chamber, which allows interpolation of the deposition rates from each source, at any position across the sample. We investigate several acquisition functions for the ML procedure. A fully Bayesian GP with the acquisition function BALM (Bayesian active learning MacKay) achieved the best performance, allowing the deposition rates for a single source to be learnt in as few as 10 experiments. Prediction accuracy for co-sputtering composition distributions was verified by comparing the SDL predictions to externally measured composition maps. Our framework dramatically increases experimental throughput by avoiding the need for time-consuming characterisation or external calibration of composition, thus demonstrating the potential of ML-guided SDLs to accelerate materials exploration, compared to the traditional Edisonian approach.
@@
Self-driving Lab, PVD, Magnetron Sputtering, Bayesian Optimisation, Gaussian processes, Active Learning, combinational thin films
@@ 1 Introduction
Thin film materials are a useful medium for high-throughput materials discovery in SDL contexts or otherwise. So far, most examples of thin film SDLs use solution coating methods for film fabrication, and some highly advanced systems exist. Key examples include systems aimed at the development of functional organic and perovskite solar cells and their constituent layers [44, 21]. In contrast, SDLs applying vacuum-based methods, Physical Vapour Deposition (PVD) or Chemical Vapour Deposition, are far less developed. This is presumably due to additional challenges of safe, high-throughput hardware automation, lack of standardised components, and greater costs involved compared to solution processing. Published examples of PVD-based SDLs include a system for optimising the reflectivity and absorptivity of sputtered single-element films [45] and another for minimising the resistance of reactively sputtered Nb-doped TiO2 deposited from individual compound targets[33]. Harris et al [12] used pulsed laser deposition to grow of WSe2 using in-situ Raman spectroscopy to estimate the growth parameters needed to maximize the film crystallinity. Further examples include Wakabayashi et al [37, 38] and Ohkubo et al [26], where SrRuO3 respectively TiN thin films were prepared using molecular beam epitaxy to obtain desired resistivity or magnetic anisotropy, or optimizing crystallinity, based on in-situ reflection high-energy electron diffraction. We note that all these examples used single material sources, with one stage of synthesis. An example using two sources was provided by Fébba et al [8], where Bayesian optimisation was used to predict composition of Znx-Ti1-x-N films deposited by reactive sputtering from Zn and Ti targets. In order to tune the power for each target, they performed in-situ plasma characterisation by optical emission spectroscopy. This method is widely applicable, although it does require external calibration.
@@
The basis for inference of composition gradients across the combinatorial thin films is characterisation of the sputter flux distribution arising from each magnetron sputtering source. The sources emit material in a directed ‘plume’ that spreads into the chamber, with an intensity and distribution that depends on the power applied to the magnetron and the chamber plasma pressure. Therefore, these variables will influence the deposition rate over the sample and the resulting composition and composition gradients on a static sample during deposition. In principle, the flux distribution might be predicted based on physical models of the sputtering process, such as developed by Ekpe et al [6]. However, speaking to the complexity of the processes involved, such models do not provide exceptionally good fits to experimental data. Simpler, geometrical models have been described that better match experimental results, based on several adjustable parameters, though without any explicit connection to physical variables such as chamber pressure or applied power [22].
For measuring sputter flux in-situ, it is common to use quartz-crystal microbalance (QCM) sensors, which monitor changes in resonant frequency of a thin quartz crystal as material is deposited upon it, to determine the mass accumulation upon them. Multiple, and even movable sensors, can also be used to perform 3D profiling of the sputter flux plume by sweeping a set of three QCMs across the deposition zone [36]. In our SDL we situate three QCMs at fixed locations around the sample, and use a geometrical flux model to interpolate the deposition rate to at other locations – specifically the ${(x,y)}$ coordinates across the area of the substrate. While QCMs are highly accurate, they only measure the total mass-deposition rate, without distinguishing contributions from different sources. Thus it is necessary to characterise the fluxes from individual sources first, before combining the results to predict the outcome of multi-source sputtering. Flux characterisation from single sources is achieved by running a sequential learning process in the SDL, training a machine learning model, in this study a Gaussian process (GP), to predict the QCM readings across the entire power-pressure parameter space for the given source. Different acquisition functions were investigated, to minimise the number of experiments needed to train for each source. This is critical to minimise source material wastage and to be able to train and retrain models as the targets are consumed over time. Trained models for several sources are then combined to predict sample composition maps, which are validated against ex-situ characterisation of samples produced by the SDL, using RBS. See Figure 1 for an overview of the workflow in the SDL.
@@ 2.2 Sputter flux model
where $\rho$ is the quartz density, $\mu$ is the shear modulus for AT-cut quartz, $f$ is the frequency of the unloaded quartz crystal, and $f_{c}$ is the measured frequency. To infer the sputter flux $\dot{m}_{x,y,z}$ at other coordinates $(x,y,z)$ between the QCM sensors, a geometric flux model according to [39] is used. In which the distribution of sputtered atoms falls off as approximately $\cos^{n}\phi$ where $\phi$ is the angle from the sputter-source surface normal shown in Figure 2.
@@
$\ \dot{m}=\frac{a(n+1)\cos\theta\cos^{n}\phi}{2\pi r^{2}},$ (2)
where $a$ is the emission rate from the source, $n$ is the order of the cosine distribution describing the shape of the flux, $r$ is the distance from the centre of the magnetron target surface to the point of deposition, and $\theta$ is the incident angle between the normal of the receiving surface (QCM or sample) and the vector from the magnetron centre, see Figure 2. For a given sputter process determined by setpoints of $P_{k}$ and $p_{\text{Ar}}$, $n$ and $a$ values, are fitted by minimising the root mean square difference between the experimental readings $\dot{m}_{i}$ at the three sensors and the flux model predictions at the corresponding QCM positions. This procedure is detailed in [34]. These calculations are implemented in the SDL code and performed automatically during operation, resulting in characterisation of the specific flux distribution for each sputtering condition tested.
@@ 2.3 Active learning framework
Figure 3: The active learning cycle begins with training models on existing data, then evaluate the acquisition function and query the new point of experiment, append the new point to the training set and repeat until the budget is spent. In each iteration, the composition map on the substrate is also obtained by using the flux model described in Section 2.2.
@@
In each active learning experiment, the model was initialised by an initial set of five points, either randomly or deliberately chosen, with another 50 being sequentially acquired during active learning. The BO parts of the workflow were implemented in Botorch [3], which uses GPyTorch [9] as the computational backend for the GP computations.
@@
To evaluate the model performance, we calculate the root mean square error (RMSE) at each iteration. RMSE is given by
$E_{\text{RMS}}=\sqrt{\frac{\sum_{i=1}^{N}||y_{i}-\hat{y}_{i}||^{2}}{N}}$ (3)
@@ 2.4 Gaussian processes
In this work, we compare a standard GP with a fully-Bayesian one. In both cases we use a Matérn 5/2 kernel with constant mean. In the fully Bayesian case, after studying the convergence, we used 32 warm-up steps and took 1024 MC samples without any thinning. To model feature importance, Automatic Relevance Determination is used, meaning each input dimension is individually scaled by lengthscales $\ell_{i}$ [40]. In $D$ dimensions the distance between points ${\bm{x}}$ and ${\bm{x}}^{\prime}$ is computed as $r^{2}=\sum_{i=1}^{D}(x_{i}-x^{\prime}_{i})^{2}/\ell_{i}^{2}$. This gives us the trainable parameters ${\bm{\theta}}=\{\mathbf{\ell},\sigma_{n},\sigma_{f},c\}$, with $\sigma_{f}$ being the signal variance and $c$ the constant mean. The fully Bayesian model uses a sparse axis-aligned subspaces prior [7], inducing a sparse structure in the inverse squared length scales $\rho_{i}\sim\mathcal{HC}(\tau)$, where $\mathcal{HC}$ denotes the half-Cauchy distribution and $\tau\sim\mathcal{HC}(\alpha)$ with $\alpha>0$ being a scalar parameter controlling the level of shrinkage. The kernel variance $\sigma_{k}^{2}\sim\mathcal{LN}(0,10^{2})$ is a log-Normal distribution.
@@ 3.1 Comparison of acquisition functions
[…]
Figure 4: ‘Ground truth’ dataset of 225 data points collected while sputtering from a Zr target, where $\dot{m}_{i}$ is the mass deposition rate measured by sensor $i$.
@@
[…]
@@
Figure 5: Comparison of various acquisition functions (a) Random query, (b) NIPV, (c) BALM and (d) BALD, for different choice of initial training points in Cases 1-3. The shaded region represents ±$\pm 1$ std. We see that BALM in Case 3 achieves the lowest $E_{\text{RMS}}$ score in the fewest queries
In the baseline case using random queries, shown in Figure 5(a), we see that the RMSE decreases as more points are acquired until the budget is spent without reaching a stable value for any Case. Case 3 achieves the lowest $E_{\text{RMS}}$ ($<2$). The variance is large and fairly constant throughout for all Cases. The large variance arises from random selection of points does not guarantee sampling from all informative parts of the space, leading to larger errors. The same effect is responsible for the steady reduction in improvement rate of $E_{\text{RMS}}$, which stays well above 0 even after 50 queries.
@@
[…]
@@ 3.2 Application and validation for co-sputtering
The starting point is for the SDL to learn the QCM deposition rates $\dot{m}_{1,2,3}$ for Cu and Sn targets individually. For the practical reasons described, we use the NIPV acquisition function and the ‘Case 3’ setup, with 30 iterations. Results of GP training are shown in 9(a), for Sn and 9(b), for Cu. As expected, the trained GPs provide excellent fits to the measured data. The different positions of the targets with respect to the QCMs account for the trends of deposition rate between the three sensors, with the nearest sensor receiving highest flux. The Sn target has a smaller range of applied power due to the lower melting point of the material.
@@
Moving from single-target sputtering to a co-sputtering situation, a few factors must be borne in mind. First, while $P_{Cu}$ and $P_{Sn}$ are independent, the $p_{Ar}$ is now a common parameter in co-sputtering. Second, the QCM sensor readings measure total mass without distinguishing the Cu and Sn contributions. Finally, it is possible that the co-sputtered elements influence each other’s deposition rates, due to chemical interactions or resputtering effects. The SDL can readily test for such effects by performing a set of co-sputtering runs with wide-ranging power and pressure conditions, and comparing the actual QCM mass-deposition rates with the sum of the predicted deposition rates from the individually-trained source GPs. In the absence of interactions, which is not unusual for metal sputtering at room temperature, the actual and measured deposition rates will be equal. This is exactly what can be seen in the example data, for Cu and Sn co-sputtering, presented in 9.
@@
Having learned the deposition rates at the QCM positions for each source, the SDL is requested to derive sputtering conditions that provide a specific composition at the centre of the substrate. To achieve this, it applies the geometrical sputter flux model described in Section 2.2, which enables interpolation of deposition rate from the QCM array to all $(x,y)$ locations across the substrate surface. This is initially done for all $P_{k}$, $p_{Ar}$ combinations, and by converting from mass- to molar-deposition rates, we derive a composition map prediction for any given combination of process conditions. By assuming bulk densities of the constituent elements and integrating over a particular deposition time, one can also predict a thickness map of the deposited film. Subsequently, the subset of conditions that result in a desired centre-point composition and film thickness can be identified. An example is shown in 9(c) for a targeted composition Cu/(Cu+Sn) = 0.66. As a result of these steps the SDL has narrowed a process parameter space containing, in this example, 875 000 possible sputtering recipes, to a few hundred recipes that provide a desired composition and film thickness. Notably, all of this was achieved automatically in the SDL based on a total of 30 minutes sputtering time (15 minutes to train GPs for each target) without producing a single physical sample.
@@
Figure 10: (a) Example composition map predicted for a co-sputtering process, expressed as ratio of Cu/(Cu+Sn) atomic densities, over a 50mm diameter wafer, showing the locations for RBS measurements on the base plane. (b) Example of RBS measurement and fitting at one measured location on a co-sputtered Cu-Sn sample. Comparison of modelled (solid bars) and measured (wireframe bars) composition at 9 positions across co-sputtered Cu-Sn samples made with different process conditions: (c) Cu: 99W, Sn: 69W, 6 mTorr, (d) Cu: 90W, Sn: 70W, 30 mTorr and (e) Cu: 51W, Sn: 33W, 6 mTorr.
@@
To validate the approach, we produce thin film Cu-Sn samples in the SDL, in this instance using three different co-sputtering conditions as indicated by the highlighted points in 9(c). Each condition should provide a centre-point composition Cu/(Cu+Sn) = 0.66 but have different composition spread, due to the different power and pressure combinations used. An example composition map prediction is shown in 10(a). RBS analysis was performed at 9 points across each sample, which were deposited onto 5 cm diameter fused silica wafers, as indicated in 10(a). Fitting of the data from each point, exemplified in 10(b), provides quantitative composition measurements allowing characterisation of the composition distribution over the wafer. The RBS results are finally compared to the predicted composition maps for each sample in 10(c)-10(e). The agreement between the predictions and the measurements is remarkable, even as the composition spread varies subtly in response to the power and pressure conditions. Indeed, the model predictions are mostly within the minor uncertainty of the RBS measurement.

### v2

@@ Abstract
Self-driving labs (SDLs) employing automation and machine learning (ML) offer great promise for accelerating materials discovery and optimisation. However, in thin film science, SDLs are mainly restricted to solution-based methods which are easier to automate, restricting access to the broader chemical space of inorganic materials. This work advances an SDL based on magnetron co-sputtering, addressing a key challenge: rapidly generating accurate composition maps of multi-element, compositionally graded thin films. Traditional ex-situ methods are slow and error-prone; instead, we present a fast, calibration-free, in-situ ML approach to predict the deposition rate using quartz-crystal microbalance (QCM) sensors. For each sputtering source, deposition rates are sequentially learned as a function of pressure and power via active learning with Gaussian processes (GPs). The final GPs are combined with a geometric flux model to interpolate deposition rates across the sample. Among several acquisition functions with random query as the baseline, the Bayesian active learning MacKay (BALM) approach yielded the best performance, requiring as few as 10 experiments per source. The model predictions for co-sputtering composition distributions were validated against external composition measurements. This framework significantly increases throughput in combinatorial sputtering studies and highlights the potential of ML-guided SDLs to surpass traditional Edisonian methods.
@@
Self-driving Lab , PVD , Bayesian Optimisation , Gaussian processes , Active Learning , combinational thin films
journal: Materials & Design
[…]
[label1]organization=Materials Science and Engineering, Uppsala University, addressline=Regementsvägen 1, city=Uppsala, postcode=75237, state=, country=Sweden
[…]
[label2]organization=Computer Science, RISE Research Institutes of Sweden, addressline=Isafjordsgatan 22, city=Kista, postcode=16440, state=, country=Sweden
[…]
[label3]organization=Information Technology, Uppsala University, addressline=Regementsvägen 1, city=Uppsala, postcode=75237, state=, country=Sweden
[…]
[label4]organization=Physics and Astronomy, Uppsala University, addressline=Regementsvägen 1, city=Uppsala, postcode=75237, state=, country=Sweden
@@ 1 Introduction
Thin film materials are a useful medium for high-throughput materials discovery in SDL contexts or otherwise. So far, most examples of thin film SDLs use solution coating methods for film fabrication, and some highly advanced systems exist. Key examples include systems aimed at the development of functional organic and perovskite solar cells and their constituent layers [14, 15]. In contrast, SDLs applying vacuum-based methods, Physical Vapour Deposition (PVD) or Chemical Vapour Deposition, are far less developed. This is presumably due to additional challenges of safe, high-throughput hardware automation, lack of standardised components, and greater costs involved compared to solution processing. Published examples of PVD-based SDLs include a system for optimising the reflectivity and absorptivity of sputtered single-element films [16] and another for minimising the resistance of reactively sputtered Nb-doped TiO2 deposited from individual compound targets[17]. [18] used pulsed laser deposition to grow of WSe2 using in-situ Raman spectroscopy to estimate the growth parameters needed to maximize the film crystallinity. Further examples include [19, 20] and [21], where SrRuO3 respectively TiN thin films were prepared using molecular beam epitaxy to obtain desired resistivity or magnetic anisotropy, or optimizing crystallinity, based on in-situ reflection high-energy electron diffraction. We note that all these examples used single material sources, with one stage of synthesis. An example using two sources was provided by [22], where Bayesian optimisation was used to predict composition of Znx-Ti1-x-N films deposited by reactive sputtering from Zn and Ti targets. In order to tune the power for each target, they performed in-situ plasma characterisation by optical emission spectroscopy. This method is widely applicable, although it does require external calibration.
@@
The basis for inference of composition gradients across the combinatorial thin films is characterisation of the sputter flux distribution arising from each magnetron sputtering source. The sources emit material in a directed ‘plume’ that spreads into the chamber, with an intensity and distribution that depends on the power applied to the magnetron and the chamber plasma pressure. Therefore, these variables will influence the deposition rate over the sample and the resulting composition and composition gradients on a static sample during deposition. In principle, the flux distribution might be predicted based on physical models of the sputtering process, such as developed by [26]. However, speaking to the complexity of the processes involved, such models do not provide exceptionally good fits to experimental data. Simpler, geometrical models have been described that better match experimental results, based on several adjustable parameters, though without any explicit connection to physical variables such as chamber pressure or applied power [27].
For measuring sputter flux in-situ, it is common to use quartz-crystal microbalance (QCM) sensors, which monitor changes in resonant frequency of a thin quartz crystal as material is deposited upon it, to determine the mass accumulation upon them. Multiple, and even movable sensors, can also be used to perform 3D profiling of the sputter flux plume by sweeping a set of three QCMs across the deposition zone [28]. In our SDL we situate three QCMs at fixed locations around the sample, and use a geometrical flux model to interpolate the deposition rate to at other locations – specifically the (x,y){(x,y)} coordinates across the area of the substrate. While QCMs are highly accurate, they only measure the total mass-deposition rate, without distinguishing contributions from different sources. Thus it is necessary to characterise the fluxes from individual sources first, before combining the results to predict the outcome of multi-source sputtering. Flux characterisation from single sources is achieved by running a sequential learning process in the SDL, training a machine learning model, in this study a Gaussian process (GP), to predict the QCM readings across the entire power-pressure parameter space for the given source. Different acquisition functions were investigated, to minimise the number of experiments needed to train for each source. This is critical to minimise source material wastage and to be able to train and retrain models as the targets are consumed over time. Trained models for several sources are then combined to predict sample composition maps, which are validated against ex-situ characterisation of samples produced by the SDL, using RBS. See Figure 1 for an overview of the workflow in the SDL.
@@ 2.2 Sputter flux model
where $\rho$ is the quartz density, $\mu$ is the shear modulus for AT-cut quartz, $f$ is the frequency of the unloaded quartz crystal, and $f_{c}$ is the measured frequency. To stay well within the applicability range of the Sauerbrey equation - where the deposited mass should be small compared to the mass of the crystal - we replaced the quartz crystals once their frequency dropped below 97.5% of the original value, 6MHz. The QCMs are water cooled to maintain constant temperature, with polished gold coating. The error in $\dot{m}_{i}$ resulting from the measurement and application of the Sauerbrey equation is approximately 5%.
To infer the sputter flux $\dot{m}_{x,y,z}$ at other coordinates $(x,y,z)$ between the QCM sensors, a geometric flux model according to [30] is used. In which the distribution of sputtered atoms falls off as approximately $\cos^{n}\varphi$ where $\varphi$ is the angle from the sputter-source surface normal shown in Figure 2.
@@
$\dot{m}=\frac{a(n+1)\cos\theta\cos^{n}\varphi}{2\pi r^{2}},$ (2)
where $a$ is the emission rate from the source, $n$ is the order of the cosine distribution describing the shape of the flux, $r$ is the distance from the centre of the magnetron target surface to the point of deposition, and $\theta$ is the incident angle between the normal of the receiving surface (QCM or sample) and the vector from the magnetron centre, see Figure 2. Values for $\theta$, $\varphi$ and $r$ are computed based on careful measurements of the chamber’s internal geometry. At the substrate centre, typical values are $\theta$ = 43.2°, $\varphi$ = 28.4°and $r$ = 155 mm. For a given sputter process determined by setpoints of $P_{k}$ and $p_{\text{Ar}}$, $n$ and $a$ values are fitted by minimising the root mean square difference between the experimental readings $\dot{m}_{i}$ at the three sensors and the flux model predictions at the corresponding QCM positions. This procedure is detailed in [31]. These calculations are implemented in the SDL code and performed automatically during operation, resulting in characterisation of the specific flux distribution for each sputtering condition tested.
@@ 2.3 Active learning framework
Figure 3: The active learning cycle begins with training three models (one for each QCM sensor mass rate) on several initial experiments, then iteratively evaluating the acquisition function and querying the next experimental point, appending the new point to the training set and repeating until the budget is spent. After training, the three sensor models are used to fit a flux distribution model (described in Section 2.2) for the source. Subsequently, multiple such flux models can be combined to predict composition maps for co-sputtering depositions.
@@
In each active learning experiment, the model was initialised by an initial set of five points, either randomly or deliberately chosen, with another 50 being sequentially acquired during active learning. The BO parts of the workflow were implemented in Botorch [32], which uses GPyTorch [33] as the computational backend for the GP computations. Our code and training dataset is available on GitHub  [^2] https://github.com/jarlsanna/gps-for-magnetron-sputtering#.
@@
For each target, single-source experiments were performed to train individual models, following the procedure in Figure 3. In real time, we only learn the deposition rate $\dot{m}$ in equation (2) from the QCM sensor located closest to the target, since this gives the strongest and most stable signal. From a physical perspective the deposition rate measured by all sensors should exhibit similar behaviour due to a common flux source. Therefore, the data points acquired for one sensor can also inform models for the others enabling rate predictions. Note that the other two models are not used during active learning. Since the sputter flux model described in Section 2.2 needs the deposition rate from all sensors, the additional models are used to obtain deposition rate predictions, which could be done offline after all experiments are performed.
To evaluate the model performance, we calculate the root mean square error ($E_{\text{RMS}}$) at each iteration. $E_{\text{RMS}}$ is given by
@@ 2.4 Gaussian processes
In this work, we compare a standard GP with a fully-Bayesian one. In both cases we use a Matérn 5/2 kernel with constant mean. An rbf kernel was also evaluated which gave similar results, but the Matérn kernel was chosen to stay more flexible. In the fully Bayesian case, after studying the convergence, we used 32 warm-up steps and took 1024 MC samples without any thinning. To model feature importance, Automatic Relevance Determination is used, meaning each input dimension is individually scaled by lengthscales $\ell_{i}$ [35]. In $D$ dimensions the distance between points $\bm{x}$} and $\bm{x}$$^{\prime}$ is computed as $r^{2}=\sum_{i=1}^{D}(x_{i}-x^{\prime}_{i})^{2}/\ell_{i}^{2}$. This gives us the trainable parameters ${\bm{\theta}}=\{\mathbf{\ell},\sigma_{n},\sigma_{f},c\}$, with $\sigma_{f}$ being the signal variance and $c$ the constant mean. The fully Bayesian model uses a sparse axis-aligned subspaces prior [36], inducing a sparse structure in the inverse squared length scales $\rho_{i}\sim\mathcal{HC}(\tau)$, where $\mathcal{HC}$ denotes the half-Cauchy distribution and $\tau\sim\mathcal{HC}(\alpha)$ with $\alpha>0$ being a scalar parameter controlling the level of shrinkage. The kernel variance $\sigma_{k}^{2}\sim\mathcal{LN}(0,10^{2})$ is a log-Normal distribution.
@@ 3.1 Comparison of acquisition functions
In this section, we present results comparing the performance of normal and fully Bayesian GPs initialised using the three different Cases described previously, using NIPV, BALM and BALD query strategies, having random experiment selection as a baseline. Concretely, performance is measured in terms of the rate of reduction of $E_{\text{RMS}}$ in equation (3), and the enhancement factor $F$ in equation (4) relative to random selection, with each queried data point. In other words, high performance corresponds to a smaller number of experiments being needed to train an accurate model.
Figure 4: ‘Ground truth’ dataset of 225 data points collected while sputtering from a Zr target, where $\dot{m}_{i}$ is the mass deposition rate measured by sensor ii.
@@
Figure 5 shows the performance of the three GP initialisation (Cases 1–3) using the different acquisition functions, in terms of $E_{\text{RMS}}$ defined by equation (3), running on the Zr dataset in Figure 4. The shaded region around each trace is the standard deviation over 10 runs with 50 queries each.
@@
Figure 5: Comparison of various acquisition functions running on the Zr dataset in Figure 4 (a) Random query, (b) NIPV, (c) BALM and (d) BALD, for different choice of initial training points in Cases 1-3. The shaded region represents $\pm 1$ std. We see that BALM in Case 3 achieves the lowest $E_{\text{RMS}}[ngcm^{-2}s^{-1}]$ score in the fewest queries
In the baseline case using random queries, shown in Figure 5(a), we see that the $E_{\text{RMS}}$ decreases as more points are acquired until the budget is spent without reaching a stable value for any Case. Case 3 achieves the lowest $E_{\text{RMS}}$ (<2<2). The variance is large and fairly constant throughout for all Cases. The large variance arises from random selection of points does not guarantee sampling from all informative parts of the space, leading to larger errors. The same effect is responsible for the steady reduction in improvement rate of $E_{\text{RMS}}$, which stays well above 0 even after 50 queries.
@@
Figure 6 shows the enhancement factor defined by equation (4) for the different acquisition functions in Cases 2 and 3, where the grey dashed line is the baseline (random sampling). We see that in Case 2, NIPV and BALM perform on average twice as well as random querying, meaning the $E_{\text{RMS}}$ of their model fits is half that of random querying, after 10 experiments. For Case 3, all query methods are better compared to random queries, with NIPV and especially BALM producing model fits nearly twice as good as random for the same number of queries. BALD is also better than random, compared to Case 2. Again, this stems back to the fact that fully Bayesian GPs are more robust, and it demonstrates the importance that the choice of initial points has for accurate model fitting; if it is possible to make a physically-relevant and unbiased choice of initialisation points, then this can be very advantageous. It is evident that BALM outperforms the others, obtaining the lowest $E_{\text{RMS}}$ in the fewest queries <10<10 with the lowest variance.
@@ 3.2 Application and validation for co-sputtering
The starting point is for the SDL to learn the QCM deposition rates $\dot{m}_{1,2,3}$ for Cu and Sn targets individually. For the practical reasons described, we use the NIPV acquisition function and the ‘Case 3’ setup, with 30 iterations. Results of GP training are shown in Figure 9(a), for Sn and Figure 9(b), for Cu. As expected, the trained GPs provide excellent fits to the measured data. The different positions of the targets with respect to the QCMs account for the trends of deposition rate between the three sensors, with the nearest sensor receiving highest flux. The Sn target has a smaller range of applied power due to the lower melting point of the material.
@@
Moving from single-target sputtering to a co-sputtering situation, a few factors must be borne in mind. First, while $P_{Cu}$ and $P_{Sn}$ are independent, the $p_{Ar}$ is now a common parameter in co-sputtering. Second, the QCM sensor readings measure total mass without distinguishing the Cu and Sn contributions. Finally, it is possible that the co-sputtered elements influence each other’s deposition rates, due to chemical interactions or resputtering effects. The SDL can readily test for such effects by performing a set of co-sputtering runs with wide-ranging power and pressure conditions, and comparing the actual QCM mass-deposition rates with the sum of the predicted deposition rates from the individually-trained source GPs. In the absence of interactions, which is not unusual for metal sputtering at room temperature, the sum of predicted deposition rates will be equal to the actual measured rates. This is exactly what can be seen in the example data, for Cu and Sn co-sputtering, presented in the Supplementary Information in Figure S2.
@@
Having learned the deposition rates at the QCM positions for each source, the SDL is requested to derive sputtering conditions that provide a specific composition at the centre of the substrate. To achieve this, it applies the geometrical sputter flux model described in Section 2.2, which enables interpolation of deposition rate from the QCM array to all $(x,y)$ locations across the substrate surface. This is initially done for all $P_{k}$, $p_{Ar}$ combinations, and by converting from mass- to molar-deposition rates, we derive a composition map prediction for any given combination of process conditions. By assuming bulk densities of the constituent elements and integrating over a particular deposition time, one can also predict a thickness map of the deposited film. Subsequently, the subset of conditions that result in a desired centre-point composition and film thickness can be identified. An example is shown in Figure 9(c) for a targeted composition Cu/(Cu+Sn) = 0.66. As a result of these steps the SDL has narrowed a process parameter space containing, in this example, 875 000 possible sputtering recipes, to a few hundred recipes that provide a desired composition and film thickness. Notably, all of this was achieved automatically in the SDL based on a total of 30 minutes sputtering time (15 minutes to train GPs for each target) without producing a single physical sample.
@@
Figure 10: (a) Example composition map predicted for a co-sputtering process, expressed as ratio of Cu/(Cu+Sn) atomic densities, over a 50mm diameter wafer, showing the locations for RBS measurements on the base plane. (b) Example of RBS measurement and fitting at one measured location on a co-sputtered Cu-Sn sample.
Figure 11: Comparison of modelled (solid bars) and measured (wireframe bars) composition at 9 positions across co-sputtered Cu-Sn samples made with different process conditions, stated in each figure.
@@
To validate the approach, we produce thin film Cu-Sn samples in the SDL, in this instance using three different co-sputtering conditions as indicated by the highlighted points in Figure 9(c). Each condition should provide a centre-point composition Cu/(Cu+Sn) = 0.66 but have different composition spread, due to the different power and pressure combinations used. An example composition map prediction is shown in Figure 10(a). RBS analysis was performed at 9 points across each sample, which were deposited onto 5 cm diameter fused silica wafers, as indicated in Figure 10(a). Fitting of the data from each point, exemplified in Figure 10(b), provides quantitative composition measurements allowing characterisation of the composition distribution over the wafer. The RBS results are finally compared to the predicted composition maps for each sample in Figure 11. The agreement between the predictions and the measurements is remarkable, even as the composition spread varies subtly in response to the power and pressure conditions. Root mean square deviations between the modelled and measured points in each 9-point composition map were 2.3%, 2% and 1.5% respectively for the examples shown in Figure 11. The centre-point compositions were 0.67. 0.65 and 0.65, compared to the targeted value of 0.66. These deviations are within the error of the RBS measurement - with the most significant error being the positioning of the sample.